OpenStudy (anonymous):

What is the simplified form of 3a^9 / 18a^3 ?

OpenStudy (anonymous):

Simplify. 3/18=1/6 and a^9/a^3=9/3=3/1= a^3/1 Therefore, you get a^3/6